OpenStudy (oleg3321):

PLEASE HELP ME. I REALLY NEED HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (oleg3321):

Create your own equation written in standard form. Determine any point that is a solution and will be on this line. Justify your answer algebraically. Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs. Rewrite the equation using function notation. Graph the function and show the intercepts.

OpenStudy (oleg3321):

@camerondoherty @phi

OpenStudy (help!!!!):

So do you know what the standard form is?

OpenStudy (oleg3321):

yes

OpenStudy (help!!!!):

make one up

OpenStudy (oleg3321):

3x + 3y = 6

OpenStudy (help!!!!):

cool now I like the y-intercept so change it into that

OpenStudy (help!!!!):

Do you know y-intercept form?

OpenStudy (oleg3321):

part a is 3x + 3y = 6 part b is (3, -1) part c ?????????? dont know

OpenStudy (oleg3321):

actually c is x-intercept = (2 , 0) y-intercept = (0 , 2) . right?

OpenStudy (help!!!!):

yup

OpenStudy (help!!!!):

third step wants you to put it into a function form. f(x) y=mx+b f(y) y=mx+b. You got it?